Hi how do you find the center and radius of a circle?
The equation is:
x^2 + y^2 = 4
Thanks

Answer by jim_thompson5910(35256)

x^2 + y^2 = 4

is in the form

(x-h)^2 + (y-k)^2 = r^2

where in this case, h = 0, k = 0 and r = 2

So the center is (h,k) = (0,0) and the radius is r = 2

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------